### v2

# Properties of neutron stars and strangeness-mixed stars from a pion mean-field approach

Abstract
We investigate the properties of the static neutron stars and strangeness-mixed stars, based on the equations of state derived from a pion mean-field approach. Using the empirical data on the pion-nucleus scattering and bulk properties of nuclear matter, we have already fixed all the parameters in a previous work, where the nucleons and hyperons were shown to be modified in various nuclear medium. In the current work, we first examine the energy and pressure inside a neutron star. We show that the central densities in various neutron stars vary within the range of $(3-4)\rho_{0}$, where $\rho_{0}$ is the normal nuclear matter density. The mass-radius relations are obtained and discussed. As the slope parameter for neutron matter increases, the radii of the neutron stars increase with their masses fixed. We also study the strangeness-mixed stars or the hyperon stars using the same sets of the parameters. As the strangeness content of strange matter increases, the binding energy per nucleon is saturated and the corresponding equation of state becomes softened. Consequently, the central densities of the strangeness-mixed stars increase. Assuming that recently observed neutron stars are the strangeness-mixed ones, we find that the central densities increase.
neutron stars, neutron matter, strange matter, pion mean-field approach
preprint: INHA-NTG-02/2024
I Introduction
Neutron stars are known to form strongly interacting cold-dense neutron matter, so that they provide a test ground for the equation of state (EoS) derived from many theories and models [1, 2]. The data on the masses and radii of the neutron stars have been compiled from astronomical observations such as low mass X-ray binaries [3, 4, 5, 6, 7, 8, 9, 10] for decades. Demorest et al. [11] observed the binary millisecond pulsar J1614-2230 from which its mass was estimated to be $(1.97\pm 0.04)M_{\odot}$. Antoniadis et al. [12] made a radio-timing observations of the PSR J0348+0432 and evaluated the mass of the neutron star as $(2.01\pm 0.04)M_{\odot}$. Fonseca et al. [13] analyzed 24 binary radio pulsars for nine years and found the mass range of the neutron stars between $1.18\,M_{\odot}$ and $1.93\,M_{\odot}$. In a recent paper [14], the pulsar mass was estimated to be $2.08\,M_{\odot}$ from the observations of PSR J0740+6620. Very recently, the observation of gravitational waves have further sharpened the data on the neutron stars [15, 16, 17]. The proposed Einstein Telescope in Europe  [18, 19, 20] and the Cosmic Explorer in the US [21, 22], which are ten times sensitive future third-generation gravitational-wave observatories, are expected to provide more precise information on the nature of neutron stars in the future. The Neutron Star Interior Composition Explorer (NICER) experiment observed PSR J0030+0451, from which the radius and mass of the neutron star were inferred to be $R=13.02_{-1.06}^{+1.24}$ km and $M=1.44_{-0.14}^{+0.15}\,M_{\odot}$ [23]. On the other hand, Riley et al. [24] estimated them as $12.39_{-0.98}^{+1,30}$ km and $2.072_{-0.066}^{+0.067}\,M_{\odot}$. We anticipate that the NICER will bring more essential data [25, 26], which can further constrain the EoS for neutron star matter (see also recent reviews and references therein [27, 28]). The mass of black widow [29], which is the fastest and heaviest known galactic neutron star, was estimated to be $M=(2.35\pm 0.17)M_{\odot}$. Its observation can also constrain the EoS.
In the present work, we aim at investigating the masses and radii of the neutron stars, based on the pion mean-field approach or the chiral quark-soliton model ($\chi$QSM) [30, 31, 32]. The idea was motivated by Witten’s seminal paper [33]: In the large $N_{c}$ (the number of colors) limit the presence of $N_{c}$ valence quarks creates the pion mean fields self-consistently. Since the mesonic quantum fluctuations are suppressed in the large $N_{c}$ limit, a baryon emerges as a state of $N_{c}$ valence quarks bound by the pion mean fields. This approach was successful in describing many properties of the low-lying baryons such as mass splittings [34, 35, 36], static properties [37, 38, 39, 40, 41, 42, 43], form factors [44, 45, 46, 47, 48], parton distribution functions [49, 50], and so on. A great merit of the pion mean-field approach lies in the fact that both the light and singly heavy baryons can be explained on an equal footing. Replacing a light valence quark with a heavy quark, we find that the singly heavy baryon arises as the bound state of the $N_{c}-1$ valence quarks by the pion mean fields. A heavy quark can be regarded as a static color source in the limit of the infinitely large mass of the heavy quark [51, 52, 53, 54, 55, 56, 57, 58, 59, 60] (see also a recent review [61]).
The pion mean-field approach can also be extended to the investigation of nuclear matter and medium modification of baryons in it. While one can introduce the quark chemical potential and examine quark matter [62, 63], it is difficult to associate quark matter with nuclear matter directly. Thus, we employ a variational approach that was successfully used in the medium modified Skyrme models [64, 65, 66, 67, 68, 69]. A virtue of the variational approach is that the dynamical parameters can be fixed by considering low-energy pion-nucleus scattering and reproducing saturation and bulk properties of homogeneous nuclear matter. In fact, the general structure of the flavor SU(3) collective Hamiltonian arises from the embedding of the SU(2) soliton into SU(3) [70, 71]. Thus, the collective Hamiltonian can be constructed based on $\mathrm{SU(2)}_{T}\otimes\mathrm{U(1)}_{Y}$ symmetry [36, 35]. Thus, while we derive the collective Hamiltonian, we fix the dynamical parameters by reproducing the properties of nuclear matter. Recently, we studied the medium modification of the low-lying SU(3) baryons in four different cases, i.e. symmetric matter, asymmetric matter, neutron matter, and strange baryonic matter consistently, based on a pion mean-field approach [72]. We were able to explain the properties of nuclear matter such as the binding energy per nucleon, symmetric energy, pressure, and medium modifications of the low-lying SU(3) baryons and singly heavy baryons [73]. Thus, we want to examine the masses and radii of neutron stars, using the EoS developed in Refs. [72, 73]. This approach has two significant virtues: First, since we have already fixed all necessary dynamical parameters, we do not need to perform any additional fitting procedure to evaluate the masses and radii of neutron stars, which will indeed play a role of the touchstone for the EoS developed in the pion mean-field approach. Second, the medium modifications of the nucleon were consistently considered.
Since we assume that neutron stars have already been cooled and stationary, we use the Tolman-Oppenheimer-Volkoff (TOV) equation for stationary neutron stars. In addition, we need to consider the effects due to the presence of leptons [74, 75, 76, 77, 78, 79, 80, 81], which brings about the contribution from the protons, in particular, at the crust of neutron stars. This will come into play to restrict the EoS [82].
The current work is organized as follows: In Sec. II, we recapitulate the pion mean-field approach, focusing on the collective Hamiltonian and a medium modification of the model using the variational approach. In Sec. III, we examine the mass, energy, and pressure densities, which will be used in the TOV equation. We will also show how we can include leptonic contributions. In Sec. IV, we discuss the results for the masses and radii of the neutron stars, and then we extend the formalism to investigate the hybrid stars, employing the strange matter. In Sec. V, we draw conclusions and summarize the current work.
II Pion mean-field approach
The $\chi$QSM is based on the low-energy QCD partition function in Euclidean space
[…]
where $\psi$ and $\psi^{\dagger}$ denote the quark fields and $\pi^{a}$ stand for the pseudo-Goldstone fields. $S_{\mathrm{eff}}[\pi^{a}]$ is the effective chiral action defined by
[…]
where $M$ represents the dynamical quark mass. It is originally the momentum-dependent one, which was derived from the fermionic zero mode of individual instantons [83, 84]. For simplicity, we turn off the momentum dependence of $M$ and introduce the regularization to tame the quark fields. $U^{\gamma_{5}}$ denotes the chiral field expressed as
$U^{\gamma_{5}}(x)=\exp{\left(i\pi^{a}(x)\lambda^{a}\gamma_{5}\right)}=U(x)\frac{1+\gamma_{5}}{2}+U^{\dagger}(x)\frac{1-\gamma_{5}}{2}$ (3)
with $U(x)=e^{i\pi^{a}(x)\lambda^{a}}$. $\hat{m}$ designates the current quark mass matrix written as
$\hat{m}=\begin{pmatrix}m_{\mathrm{u}}&0&0\\ 0&m_{\mathrm{d}}&0\\ 0&0&m_{\mathrm{s}}\end{pmatrix}=m_{0}\mathbf{1}+m_{3}\lambda^{3}+m_{8}\lambda^{8}$ (4)
with
$m_{0}=\frac{m_{\mathrm{u}}+m_{\mathrm{d}}+m_{\mathrm{s}}}{3},\;\;m_{3}=\frac{m_{\mathrm{u}}-m_{\mathrm{d}}}{2},\;\;m_{8}=\frac{m_{\mathrm{u}}+m_{\mathrm{d}}-2m_{\mathrm{s}}}{2\sqrt{3}}.$ (5)
Here, $\lambda^{a}$ represent the Gell-Mann matrices for the flavor SU(3) group and $m_{\mathrm{u}}$, $m_{\mathrm{d}}$, and $m_{\mathrm{s}}$ denote the current up, down, and strange quark masses, respectively. The effective chiral action contains all orders of the effective chiral Lagrangians with the low-energy constants, which can be derived from the gradient expansion [30, 85].
The classical nucleon mass or the chiral soliton mass can be obtained by solving the nucleon correlation function. By solving the correlation function, we mean that the $N_{c}$ valence quarks are positioned in the lowest level bound by the pion mean fields produced by the presence of the $N_{c}$ valence quarks self-consistently. Since the mesonic quantum fluctuations are known to be suppressed in the large $N_{c}$ limit [33], we can integrate over the pseudo-Goldstone boson fields in Eq. (1) around the saddle point $U_{c}(\bm{x})$. The equation of motion for the quarks with the classical pion fields can be minimized by the Hartree approximation, we obtain the pion mean fields with hedgehog symmetry
$U_{c}(\bm{x})=\exp\left(i\bm{\tau}\cdot\bm{n}P(r)\right),$ (6)
where $P(r)$ is called the profile function. In the current work, we will not follow the self-consistent procedure to determine the profile functions and classical soliton mass. Instead, we will maximally use the $\mathrm{SU(2)}_{T}\otimes\mathrm{U(1)}_{Y}$ symmetry that arises from the embedding of the SU(2) chiral soliton into SU(3)
$U(\bm{x})=\begin{pmatrix}U_{c}(\bm{x})&0\\ 0&1\end{pmatrix}.$ (7)
The dynamical parameters will be determined by using the experimental and empirical data. This method is often called “a model-independent” approach for the chiral soliton, which has a virtue when it is extended to the investigation of bulk properties of nuclear matter and baryon properties in it. Since the classical soliton does not carry quantum numbers of baryons, we need to quantize it. As mentioned already, the quantum fluctuations of the pion fields are suppressed by the large $N_{c}$ approximation. On the other hand, the fluctuation of the pion field to the zero-mode directions, which are always related to translational and rotational symmetries, must be completely treated. Rotational zero modes determine the quantum numbers of baryons such as the spin and isospin. For details, we refer to Refs. [31, 32].
Having performed the rotational zero-mode quantization, we derive the collective Hamiltonian [34, 36, 35]
$H=M_{{\rm cl}}\;+\;H_{\mathrm{rot}}\;+\;H_{\mathrm{sb}}+\;H_{\mathrm{em}},$ (8)
where $M_{\mathrm{cl}}$ denotes the mass of the classical nucleon. $H_{\mathrm{rot}}$ stands for the rotational $1/N_{c}$ corrections arising from the zero-mode quantization
$H_{\mathrm{rot}}$ $=\frac{1}{2I_{1}}\sum_{i=1}^{3}\hat{J}_{i}^{2}\;+\;\frac{1}{2I_{2}}\sum_{p=4}^{7}\hat{J}_{p}^{2},$ (9)
where the $I_{1,2}$ are the moments of inertia and $\hat{J}_{a}$ ($a=1,\cdots 8$) are the generators of the $\mathrm{SU(3)}$ group, which are directly related to the right angular velocities $\Omega_{a}$. The eighth component of $J_{a}$ is constrained to be
$J_{8}=-\frac{N_{c}}{2\sqrt{3}}B=-\frac{\sqrt{3}}{2},$ (10)
[…]
The third term in Eq. (8) stands for both the isospin and SU(3) flavor symmetry breaking part
$H_{\mathrm{sb}}$ $=\left(m_{\mathrm{d}}-m_{\mathrm{u}}\right)\left(\frac{\sqrt{3}}{2}\,\alpha\,D_{38}^{(8)}(\mathcal{A})\;+\;\beta\,\hat{T_{3}}\;+\;\frac{1}{2}\,\gamma\sum_{i=1}^{3}D_{3i}^{(8)}(\mathcal{A})\,\hat{J}_{i}\right)$ (12)
$+\;\left(m_{\mathrm{s}}-\bar{m}\right)\left(\alpha\,D_{88}^{(8)}(\mathcal{A})\;+\;\beta\,\hat{Y}+\;\frac{1}{\sqrt{3}}\,\gamma\sum_{i=1}^{3}D_{8i}^{(8)}(\mathcal{A})\,\hat{J}_{i}\right),$ (13)
where $\alpha$, $\beta$, $\gamma$ are expressed in terms of the moments of inertia
$\alpha$ $=-\left(\frac{2}{3}\frac{\Sigma_{\pi N}}{m_{\mathrm{u}}+m_{\mathrm{d}}}-\frac{K_{2}}{I_{2}}\right),\;\;\;\;\beta=-\frac{K_{2}}{I_{2}},\;\;\;\;\gamma=2\left(\frac{K_{1}}{I_{1}}-\frac{K_{2}}{I_{2}}\right).$ (14)
Here $K_{1,2}$ represents the anomalous moments of inertia of the rotating soliton. The $\overline{m}$ is the average value of the up and down current quark masses, i.e. $\overline{m}=(m_{\mathrm{u}}+m_{\mathrm{d}})/2$. The $D_{ab}^{(\mathcal{R})}(\mathcal{A})$ are the SU(3) Wigner $D$ functions in a given representation $\mathcal{R}$. The explicit form of the electromagnetic (EM) self-energy term $H_{\mathrm{em}}$ is given in Ref. [35]. However, the effects of the EM self-energies are negligibly small for medium modification [65, 66], we ignore them in the current work.
The dynamical parameters $\alpha$, $\beta$, and $\gamma$ can be evaluated within the self-consistent SU(3) $\chi$QSM. However, we will adopt the model-independent approach, as previously mentioned. So, we will fix all dynamical parameters by using the experimental data rather than by computing within the model [86, 36]. Furthermore, we employ the variational method to modify the collective Hamiltonian (8) in nuclear medium. We briefly recapitulate the general formalism for the medium modification of $H$. For details, we refer to Refs. [72, 73].
Before we proceed to discuss the density dependence of the binding energy per nucleon, we will briefly mention the corrections from the translational zero modes that make the nucleon carry the linear momentum. The zero-mode translation can be included by translating the chiral soliton as
$U_{x}(\bm{x}-\bm{X}(\tau))=T_{\bm{X}(\tau)}U_{c}(\bm{x})T_{\bm{X}(\tau)}^{\dagger},$ (15)
where $\bm{X}(\tau)$ is the translational zero modes conjugate to the linear momentum of the soliton. $T_{\bm{X}(\tau)}$ is a unitary SU(2) matrix that governs the corresponding translation. Once we carry out the quantization by expanding the effective chiral action with respect to the velocity $\dot{\bm{X}(\tau)}$ to the second order, and compute the functional integration over translational zero modes, we have kinetic corrections to the collective Hamiltonian given in Eq. (8)
$H^{\prime}=H+\frac{\hat{\bm{P}}^{2}}{2M_{\mathrm{cl}}},$ (16)
where the kinetic term is of the order $\mathcal{O}(N_{c}^{0})$ in the large $N_{c}$ limit. $\hat{\bm{P}}$ is the linear momentum operator conjugate to $\hat{\bm{X}}$. However, there is one important caveat. The classical mass $M_{\mathrm{cl}}$ already contains spurious contributions from the center-of-mass motion, which must be subtracted. This process is nontrivial as was discussed in depth within the Skyrme models [87, 88], with $1/N_{c}$ pion-loop corrections included. Technically, this has resemblance with the random phase approximation in many-body theory. We may expect that two main subtraction terms expressed as
$\Delta E=-\frac{\langle P^{2}\rangle}{2M_{\mathrm{cl}}}-\frac{\langle J^{2}\rangle}{2I},$ (17)
where $\langle p^{2}\rangle$ and $\langle J^{2}\rangle$ denote the average values of $P^{2}$ and $J^{2}$. The derivation of these values is beyond the current theoretical framework. Thus, we assume that the classical mass of the nucleon effectively carry the corrections from the kinetic energy.
In Refs.  [72, 73], three different types of nuclear environments were considered, i.e. the isospin symmetric, isospin asymmetric, and strangeness-mixed nuclear matters. Thus, we introduce three corresponding parameters as follows:
$\lambda=\frac{n}{n_{0}},\quad\delta=\frac{N-Z}{A},\quad\delta_{s}=\frac{N_{s}}{A},$ (18)
where $\lambda$ is the density of nuclear matter normalized by the normal nuclear matter density $\rho_{0}\simeq 2.7\times 10^{14}\,\mathrm{g/cm^{3}}$ ($n_{0}=0.16\,\mbox{nucleons per $\mathrm{fm}^{3}$}$). The isospin asymmetry in nuclear matter is controlled by $\delta$, which is related to the number of neutrons $N$, of protons $Z$, and of baryons $A$. The last variational parameter $\delta_{s}$ is responsible for the strangeness mixing ratio that is proportional to the number of baryons $N_{s}$ with strangeness $s=|S|$. Using these three parameters, we can define the binding energy per baryon as
$\varepsilon(\lambda,\delta,\,\delta_{1},\delta_{2},\delta_{3})$ $=\Delta M_{N}\left(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}\right)\left(1-\sum_{s=1}^{3}\delta_{s}\right)+\frac{1}{2}\delta\,\Delta M_{np}\left(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}\right)$ (19)
$+\sum_{s=1}^{3}\delta_{s}\Delta M_{s}\left(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}\right).$ (20)
See Ref. [72] for details. As a result, we have the following density-dependent classical nucleon mass $M_{\mathrm{cl}}^{\ast}$, the moments of inertia $I_{\mathrm{1,2}}^{\ast}$, the in-medium modified isospin symmetry breaking [^1] Here, the strong part of the in-medium modified effects on isospin symmetry breaking are presented. $E_{\mathrm{iso}}^{\ast}$ and the $\mathrm{SU(3)}$ symmetry breaking $E_{\mathrm{str}}^{\ast}$ terms:
$M_{\mathrm{cl}}^{\ast}$ $=M_{\mathrm{cl}}f_{\mathrm{cl}}(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}),$ (21)
$I_{\mathrm{1}}^{\ast}$ $=I_{\mathrm{1}}f_{1}(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}),$ (22)
$I_{\mathrm{2}}^{\ast}$ $=I_{\mathrm{2}}f_{2}(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}),$ (23)
$E_{\mathrm{iso}}^{\ast}$ $=\left(m_{\mathrm{d}}-m_{\mathrm{u}}\right)\frac{K_{1,2}}{I_{1,2}^{*}}f_{0}(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}),$ (24)
$E_{\mathrm{str}}^{\ast}$ $=\left(m_{\mathrm{s}}-\bar{m}\right)\frac{K_{1,2}}{I_{1,2}^{*}}f_{\mathrm{s}}(\lambda,\delta,\delta_{1},\delta_{2},\delta_{3}),$ (25)
where $f_{\mathrm{cl}}$, $f_{\mathrm{0,1,2}}$, and $f_{s}$ are given as the functions of the baryon environment densities and other medium parameters. These functions are explicitly written as
$f_{\mathrm{cl}}(\lambda)$ $=\left(1+C_{\mathrm{cl}}\lambda\right),$ (26)
$f_{1,2}(\lambda)$ $=\left(1+C_{1,2}\lambda\right),$ (27)
$f_{0}(\lambda,\delta)$ $=1+\frac{C_{\mathrm{num}}\lambda\,\delta}{1+C_{\mathrm{den}}\lambda}\,,$ (28)
$f_{s}(\lambda,\delta_{s})$ $=1+g_{s}(\lambda)\delta_{s}$ (29)
$g_{s}(\lambda)$ $=sg(\lambda),$ (30)
$g(\lambda)$ $=\left(6\frac{K_{2}}{I_{2}}+\frac{K_{1}}{I_{1}}\right)^{-1}$ (31)
×$\hskip-28.45274pt\times\frac{5(M_{cl}-M_{\mathrm{cl}}^{\ast}+E_{(1,1)1/2}-E_{(1,1)1/2}^{*})}{3(m_{s}-\hat{m})},$ (32)
where the $C_{\mathrm{cl}}$, $C_{1}$, $C_{2}$, $C_{\mathrm{num}}$, $C_{\mathrm{den}}$, which are fixed by reproducing the bulk properties of nuclear matter [72]
$C_{\mathrm{cl}}=-0.0561,\quad C_{1}=0.6434,\quad C_{2}=-0.1218,\quad C_{\mathrm{num}}=65.60,\quad C_{\mathrm{den}}=0.60.$ (33)
We find that the present medium-modified EoS describes very well bulk properties of nuclear matter and symmetry energies up to density $\sim 3\rho_{0}$. Even at higher densities, the results are in reasonable agreement with empirical data. We want to examine the EoS developed in Ref. [72] by computing the mass-radius relation of neutron stars in the current work. For simplicity, we also introduce the common strangeness mixing parameter $\chi$ as
$\sum_{\mathrm{s}}\mathrm{s}\delta_{\mathrm{s}}=\chi,$ (34)
where $s$ is the strangeness of the hyperon. Then the binding energy per baryon is given as a function of three parameters $\varepsilon=\varepsilon(\lambda,\delta,\chi)$.
The nuclear symmetry energy is defined as the second derivative of the binding energy with respect to $\delta$
$\varepsilon_{\mathrm{sym}}\left(\lambda\right)$ $=\frac{1}{2!}\left.\frac{\partial^{2}\varepsilon\left(\lambda,\delta,0,0,0\right)}{\partial\delta^{2}}\right|_{\delta=0}.$ (35)
$\varepsilon_{\mathrm{sym}}$ can be expanded in the vicinity of the saturation point $\lambda=1$
$\varepsilon_{\mathrm{sym}}\left(\lambda\right)$ $=a_{\mathrm{sym}}+\frac{L_{\mathrm{sym}}}{3}\left(\lambda-1\right)+K_{\mathrm{sym}}\frac{\left(\lambda-1\right)^{2}}{18}+\cdots,$ (36)
where $a_{\mathrm{sym}}$ denotes the nuclear symmetry energy at the saturation point, $L_{\mathrm{sym}}$ stands for that of its slope parameter $L_{\mathrm{sym}}$, and $K_{\mathrm{sym}}$ represents the asymmetric part of the incompressibility. They are expressed as
$a_{\mathrm{sym}}$ $=-\frac{9}{20}\frac{C_{\mathrm{num}}\left(b-7r/18\right)}{\left(1+C_{\mathrm{den}}\right)},$ (37)
$L_{\mathrm{sym}}$ $=-\frac{27}{20}\frac{C_{\mathrm{num}}\left(b-7r/18\right)}{\left(1+C_{\mathrm{den}}\right)^{2}},$ (38)
$K_{\mathrm{sym}}$ $=\frac{81C_{\mathrm{den}}C_{\mathrm{num}}(b-7r/18)}{10(1+C_{\mathrm{den}})^{3}},$ (39)
where $b=(m_{\mathrm{d}}-m_{\mathrm{u}})\beta$ and $r=(m_{\mathrm{d}}-m_{\mathrm{u}})\gamma$. The empirical values of $a_{\mathrm{sym}}$ and $L_{\mathrm{sym}}$ are known to be $a_{\mathrm{sym}}=(31.7\pm 3.2)$ MeV and $L_{\mathrm{sym}}=(58.7\pm 28.1)$ MeV, respectively [2]. In the current work, we take four different sets of parameters for $a_{\mathrm{sym}}$ and $L_{\mathrm{sym}}$ as shown in Table 1.
Table 1: Four different sets of the parameters for the symmetry energy and slope parameter at the normal nuclear matter density $\rho_{0}$.
$a_{\rm sym}$ [ MeV] $L_{\rm sym}$ [MeV] $C_{\rm num}$ $C_{\rm den}$
Set I 30 50 69.20 0.80
Set II 32 50 78.72 0.92
Set III 30 60 57.66 0.50
Set IV 32 60 65.60 0.60
The density dependence of the symmetry energy is constrained by the experimental data on heavy-ion collision [89, 90, 91, 92]. In the previous work [72], we showed that the results for $\varepsilon(\lambda)$ are in good agreement with the APR predictions [93] and IAS constraints [94]. The values of $a_{\mathrm{sym}}$ and $L_{\mathrm{sym}}$ are also very similar to those in various theoretical works [1, 10, 95, 96]. As will be shown later, the results for the masses and radii of the neutron stars are not sensitive to changes of the symmetry energy, whereas its slope parameter has certain effects on them. In Ref. [72], we also discussed as to how the parameters, $K_{\mathrm{sym}}$ and related incompressibilities, were determined in detail. The results for both the energy and pressure indicate that the slope parameter of the symmetric energy will influence the mass-radius relations of the neutron stars.
Figure 1: The energy and pressure as functions of $\lambda$. The parameters in the legend correspond to those given in Table 1. The green area represents the results from Ref. [97].
The left panel of Fig. 1 depicts the binding energy per nucleon for neutron matter as a function of $\lambda$, of which the expression is given by Eq. (20) with $\delta=1$ and $\delta_{s}=0$. The current results are in good agreement with those obtained in chiral effective theory [97] with the three-loop order corrections and three-body interaction considered. We find one interesting feature of the slope parameter $L_{\mathrm{sym}}$. When $\lambda$ is small, the symmetry energy governs the behavior of the energy per nucleon. As $\lambda$ increases, the slope parameter takes charge of it. The right panel of Fig. 1 draws the pressure as a function $\lambda$. We also compare the results with those from Ref. [97]. As observed in the right panel of Fig. 1, a larger value of the slope parameter, i.e., $L_{\mathrm{sym}}=32$ MeV enhances the pressure. On the other hand, the change of the symmetry energy shows only a tiny effect on the pressure.
III Energy densities and pressures for neutron stars
To study the mass-radius relation of the neutron stars, we consider static and spherically symmetric non-rotating neutron stars, which satisfy the TOV equation. The mass function of the neutron star $M(r)$ is given as a function of radius $r$
$\mathcal{M}(r)=4\pi\int_{0}^{r}\mathrm{d}r\,r^{2}\mathcal{E}(r),$ (40)
where $\mathcal{E}(r)$ stands for the energy density as a function of $r$. Then the total mass of the neutron star with radius $R$ is defined by
$M=\mathcal{M}(R).$ (41)
The pressure $P(r)$ is determined by solving the TOV equation
$\frac{dP(r)}{dr}$ $=\frac{\mathcal{E}(r)\mathcal{M}(r)}{r^{2}}\left(1-\frac{2\mathcal{M}(r)}{r}\right)^{-1}\left(1+\frac{P(r)}{\mathcal{E}(r)}\right)\left(1+\frac{4\pi r^{2}P(r)}{\mathcal{M}(r)}\right)$ (42)
in natural units $c=G=1$. The boundary conditions for the mass and energy densities at the center of the neutron star are given as
$\mathcal{M}(0)=0,\quad\mathcal{E}(0)=\mathcal{E}_{\mathrm{c}},$ (43)
where the $\mathcal{E}_{\mathrm{c}}$ is the central energy density of the neutron star. The stability condition requires the local pressure to be zero at the surface of the neutron star
$P(R)=0.$ (44)
To obtain the profile of a neutron star, we need to know the energy dependence of the pressure
$P=P(\mathcal{E}),$ (45)
which can be derived from the following equation
$P(\lambda,\delta,\chi)$ $=\rho_{0}\lambda^{2}\frac{\partial\varepsilon(\lambda,\delta,\chi)}{\partial\lambda},$ (46)
$\mathcal{E}(\lambda)$ $=\left[\varepsilon(\lambda,\delta,\chi)+m_{N}\right]\lambda\rho_{0}.$ (47)
Figure 2: The energy density and pressure of the neutron star as functions of $r$.
The numerical result for $\mathcal{E}(r)$ is drawn in the left panel of Fig. 2 with the value of the incompressibility $K_{0}=240$ MeV. The results are rather stable as the parameters are changed. As expected, the energy density falls off monotonically, as $r$ increases, and then it slowly vanishes at the surface of the neutron star. The right panel of Fig. 2 shows the interior pressure of the neutron star as a function of $r$, which is obtained by solving Eq. (42). The pressure also drops off as $r$ increases, and vanishes at around 11-12 km, which corresponds to the radius of the neutron star. Using Eq. (41) and (44), we can derive the masses and radii of the neutron stars.
Although the leptons inside a neutron star do not participate in the strong interaction, this will cause certain contributions from the protons in the crust of the neutron star. We use a relativistic Fermi gas model to describe the electrons inside a the neutron star [98, 76], so that we obtain the energy density and pressure of the electron gas as
$\mathcal{\epsilon}_{e}$ $=\frac{3}{4}n{e}\mu_{e},\;\;\;P_{\mathrm{e}}=\frac{1}{4}n_{e}\mu_{e},$ (48)
where $n_{e}$ denotes the number density of the electrons, and $\mu_{e}$ is the chemical potential of the electron.
Since we consider the beta-stable condition, the chemical potential of the electron can be expressed in terms of the symmetry energy as [78]
$\mu_{e}=4\delta E_{\mathrm{sym}}(\lambda).$ (49)
Since the neutron star satisfies the charge neutrality condition, the electron number density must be the same as the proton number density, which brings about the contribution from the protons. Thus, the symmetry energy depends not only on the density of the neutron star but also on the fraction of the protons. We fix the electron and proton fractions to be $n_{e}=n_{p}=0.1$ in the current work, assuming that the leptonic effects mainly come from the crust of the neutron star. So, the electron affects only the subnuclear density region, i.e., $\lambda<1$ [78].
IV Results and discussions
Before we proceed to discuss the results, we want to emphasize that the current theoretical framework is based on linear density approximation. Moreover, we do not include quark matter in the current work. It implies that we can only describe physical observables in medium approximately up to $\rho\approx 4\rho_{0}$. If the medium density becomes larger than $4\rho_{0}$, the speed of sound would break the causality. Thus, in the forthcoming discussions, we will restrict ourselves to lower densities than $4\rho_{0}$.
IV.1 Neutron stars
Since all the parameters were determined in the previous work [72], we do not change any parameters for evaluating the masses and radii of the neutron stars. Using the four different sets of parameters listed in Table 1, we proceed to compute the mass-radius relations of the neutron stars.
Figure 3: Masses of the neutron star are drawn as functions of the dimensionless central density $\lambda=\rho/\rho_{0}$ with $\rho_{0}=0.16\,\mathrm{fm}^{-3}$ in unit of the solar mass $M_{\odot}$. The data for the orange, blue, and green bars are taken from Refs. [14, 12, 13].
In Fig. 3, we show the masses of the neutron star in units of the solar mass as functions of dimensionless central density $\lambda$ with the four different parameter sets used. While the results are not sensitive to $a_{\mathrm{sym}}$, they depend on $L_{\mathrm{sym}}$ in lower nuclear matter density. The results indicate that the central density of the neutron stars observed in Refs. [14, 12, 13] is approximately $\lambda\approx 4$ or $\rho\approx 0.64\,\mathrm{fm}^{-3}$. However, when heavy neutron stars are considered, quark matter must come into a significant role, which we did not include in the current work. This, we present the results for the masses of neutron stars up to $\lambda\approx 4$.
It is interesting to compare the current results with those from other approaches. In Ref. [93], Akmal et al. constructed the EoS by using the Argonne nucleon-nucleon potential (A18), Urbana three-nucleon interactions, the boost interaction, and the estimated maximum masses of $M=2.2M_{\odot}$ at $\lambda\approx 6$. Baym et al. [28] also obtained the maximum masses of $M=2.2M_{\odot}$ saturated at around $\lambda\approx 6$. On the other hand, Brandes et al. [99] incorporated the data on the newly observed neutron called the black-widow (PSR J0952-0607) [29], and estimated $\lambda\approx 4$ for $M=2.1\,M_{\odot}$. Vijayan et al. [100] reported very interesting results. Using two different EoS and considering the medium modification of the pion mass, they found that the central density reaches $\lambda\approx 6$ for $M=(1.7-2.3)M_{\odot}$. Since the current framework has naturally taken into account the medium modifications of hadrons, the present results are in good agreement with those from Ref. [29].
Figure 4: Mass-radius relations of neutron stars. The black curves represent the results with the four different sets of parameters given in Table 1. The value of the incompressibility is chosen as $K_{0}=240$ MeV. The red and green lines with stars are taken from the LIGO experimental data [101]. The data for the orange, blue, and green bars are taken from Ref. [14, 12, 13].
Figure 4 depicts the mass-radius relations for the neutron stars. As already discussed above, the mass-radius relations are not sensitive to $a_{\mathrm{sym}}$, the radius tends to increase as the slope parameter $L_{\mathrm{sym}}$ increases, in particular, when the neutron matter density is low. As the density increases, the dependence of the mass-radius relations on $L_{\mathrm{sym}}$ becomes smaller. Figure 4 also indicates that the central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] are estimated to be around $3.5\rho_{0}$. As mentioned previously, by including the leptonic effects, the radius of the neutron stars tend to decrease. They mainly contribute to lower-density regions.
Figure 5: Dependence of the neutron star mass-radius relation on the incompressibility, $K_{0}$. The values of the symmetry energy at saturation density and the corresponding slope parameter are chosen as $a_{\rm sym}=32$ MeV and $L_{\rm sym}=60$ MeV, respectively. The observational and empirical data are the same as in Fig. 4.
The empirical values of the nuclear matter incompressibility $K_{0}$ are given in a wide range. For example, the value of $K_{0}$ was often predicted as $K_{0}\approx(290\pm 70)$ MeV [102, 103, 104, 105, 106, 107]. In Ref. [108], however, a smaller value was extracted from the data on the isoscalar giant monopole resonance in ${}^{208}\mathrm{Pb}$, based on relativistic mean-field models: $K_{0}=(248\pm 8)$ MeV. In Ref. [109], $K_{0}=(240\pm 20)$ MeV was extracted from a reanalysis of the data on the energies of the giant monopole resonances in even-even 112-124Sn and 106,100-116Cd and earlier data on $58\leq A\leq 208$ nuclei. A similar value was also estimated in Ref. [110]. Thus, while we choose $K_{0}=240\,\mathrm{MeV}$, it is of great interest to examine the dependence of the mass-radius relations of the neutron stars on $K_{0}$. In Fig. 5, we demonstrate the mass-radius relations of the neutron stars, varying the incompressibility of nuclear matter $K_{0}$ from 210 MeV to 260 MeV. We take the values of $a_{\mathrm{sym}}=32$ MeV and $L_{\mathrm{sym}}=60$ MeV. The tendency of the mass-radius relations is obvious in general. If we increase $K_{0}$ from 220 MeV to 260 MeV, the mass-radius relations remain as almost the same shape but are shifted to the right. So, the radii of the neutron stars get larger as $K_{0}$ increases.
In Table 2, we list the results for the central densities and the radii, which correspond to two different values of the neutron star masses, i.e. $M=1.4M_{\odot}$ and $M=2.0M_{\odot}$. We take different values of $K_{0}$ with two different sets of parameters for neutron matter. The results show that as $K_{0}$ increases, the central density decreases, so the radius of the neutron star enlarges.
Table 2: central densities and radii of the neutron stars with two different values of the mass. We take three different values of the incompressibility of nuclear matter with the two different sets of symmetry energy and slope parameters at saturation density $\rho_{0}$. $n$ denotes the central density of neutron matter, $\lambda$ stands for the corresponding dimensionless central density, and $R$ designates the radius of the neutron star. Note that $n_{0}$ corresponds to $\rho_{0}=2.677\times 10^{14}$ g/cm3.
[…]
$K_{0}$, $a_{\mathrm{sym}}$, $L_{\mathrm{sym}}$ $n$ $\lambda$ $R$ $n$ $\lambda$ $R$
$[\mathrm{MeV}]$ $[\mathrm{fm}^{-3}]$ [dimensionless] [km] $[\mathrm{fm}^{-3}]$ [dimensionless] [km]
210 MeV, 30 MeV, 50 MeV 0.502 3.136 11.514 0.672 4.198 11.308
240 MeV, 30 MeV, 50 MeV 0.429 2.681 12.121 0.549 3.430 12.094
270 MeV, 30 MeV, 50 MeV 0.385 2.406 12.542 0.477 2.983 12.642
210 MeV, 32 MeV, 50 MeV 0.485 3.030 11.490 0.646 4.036 11.281
240 MeV, 32 MeV, 50 MeV 0.454 2.839 12.110 0.587 3.666 12.074
270 MeV, 32 MeV, 50 MeV 0.434 2.712 12.547 0.550 3.435 12.629
210 MeV, 30 MeV, 60 MeV 0.480 2.997 11.858 0.658 4.115 11.513
240 MeV, 30 MeV, 60 MeV 0.413 2.584 12.435 0.538 3.363 12.299
270 MeV, 30 MeV, 60 MeV 0.372 2.328 12.818 0.468 2.927 12.834
210 MeV, 32 MeV, 60 MeV 0.485 3.032 11.826 0.663 4.146 11.473
240 MeV, 32 MeV, 60 MeV 0.417 2.609 12.426 0.541 3.384 12.272
270 MeV, 32 MeV, 60 MeV 0.375 2.347 12.831 0.470 2.944 12.820
IV.2 Strageness-mixed stars
A virtue of the current approach is that we can easily incorporate the hyperons into nuclear matter. As shown in Eq. (18), the parameter $\delta_{s}$ controls the proportionality of the hyperons with strangeness $s=|S|$. Thus, the energy density can depend on $\delta_{s}$ as in Eq. (20). Once we construct the EoS for baryonic matter with strangeness, we can evaluate the mass-radius relations of hyperon stars. Before we compute them, we will recapitulate how baryonic matter can be formulated within the present approach. Note that we will not introduce any new parameters but introduce the hyperons with strangeness $s=|S|$.
We first expand the binding energy per baryon given in Eq.(20) with respect to isospin-asymmetry parameter $\delta$ and strangeness-mixing parameters $\delta_{s}$
$\varepsilon(\lambda,\delta,\delta_{1},\dots)$ $=\varepsilon_{V}\left(\lambda\right)+\varepsilon_{\mathrm{sym}}\left(\lambda\right)\delta^{2}+\sum_{s=1}^{3}\left.\frac{\partial\,\varepsilon\left(\lambda,\,\delta,\,\delta_{1},\dots\right)}{\partial\,\delta_{s}}\right|_{\delta=\delta_{1}=\dots=0}\delta_{s}$ (50)
$+\frac{1}{2}\sum_{s,p=1}^{3}\left.\frac{\partial^{2}\,\varepsilon\left(\lambda,\,\delta,\,\delta_{1},\dots\right)}{\partial\,\delta_{s}\partial\delta_{p}}\right|_{\delta=\delta_{1}=\dots=0}\delta_{s}\delta_{p}+\cdots,$ (51)
where the first and second terms denote the volume and symmetry energies for ordinary nuclear matter. It is obvious that the linear terms in $\delta$ are absent due to the quadratic dependence of the binding energy per baryon on it. Assuming that higher-order terms in $\delta_{s}$ are negligible, we choose $f_{s}$’s similar to $f_{0}$ given in Eq. (28). We will parametrize $f_{s}$ that is independent of $\delta$, so that we keep the nonstrange sector intact. For detailed analysis, we refer to Ref. [72].
Figure 6: The energy and pressure for baryonic matter as functions of $\lambda$. The sum of the parameters $\delta_{s}$ varies from 0 to 1.
In Fig. 6, we draw the results for the energy and pressure as functions of $\lambda$ with the sum of the parameters $\delta_{s}$ changed from 0 to 1. While $\sum\delta_{s}=0$ corresponds to pure neutron matter, $\sum\delta_{s}=1$ indicates pure strange matter. The left panel of Fig. 6 shows that as the number of the strange quarks increases, the energy per nucleon starts to decrease. When $\lambda$ reaches about 0.160.160.16, $\varepsilon(\lambda)$ becomes negative and turns positive, which exhibits a behavior similar to normal nuclear matter. As the content of the strangeness increases, the saturation point gets lower till it reaches 5.97-5.97- 5.97 MeV at $\lambda\simeq 1$ for pure strange matter. However, strangeness-mixed matter has a shallow saturation point in comparison with nuclear matter [72]. Consequently, the pressure of strangeness-mixed matter becomes weaker as the content of the strangeness increases, as demonstrated in the right panel of Fig. 6. Interestingly, the pressure is lessened by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softer than neutron matter, as discussed in Refs. [111, 112, 113].
Figure 7: Masses of the strangeness-mixed stars in unit of the solar mass as functions of $\lambda$ with the strangeness content varied from zero to 1. The values of the symmetry energy, the slope parameter, and the incompressibility at the saturation density are taken as $a_{\mathrm{sym}}=32\,\mathrm{MeV}$, $L_{\mathrm{sym}}=60\,\mathrm{MeV}$, and $K_{0}=240$ MeV. The observational and empirical data are the same as in Fig. 4.
In Fig. 7, we draw the results for the masses of the strangeness-mixed stars as functions of $\lambda$. The short dashed curve depicts the results for the neutron star mass with the parameters $a_{\mathrm{sym}}=32\,\mathrm{MeV}$, $L_{\mathrm{sym}}=60\,\mathrm{MeV}$, and $K_{0}=240$ MeV used. If we add the strange quarks by 25 %, the mass is not much changed except for the lower density region ($\lambda\geq 3$), as shown by the dashed curve. If we increase the strange quark content, the masses of the strangeness-mixed star become smaller at the same central density. To get the masses of the known neutron stars with higher strangeness content, the central densities should become larger. It implies that as the EoS gets softer, the central density becomes larger. Similar conclusions are found in Refs. [111, 112, 113, 114, 115, 116].
Figure 8: The mass-radius relation of a hyperon-mixed star that has the different portion of hyperons. The parameters of EOS are fixed at values $K_{0}=240$ MeV, $a_{\mathrm{sym}}=32\,\mathrm{MeV}$ and $L_{\mathrm{sym}}=60\,\mathrm{MeV}$. The portion of protons is equal to zero. The results from the experiment and analyses are given in Fig. 4.
The numerical results for the mass-radius relations of the strangeness-mixed stars are depicted in Fig. 8. We fix the parameters for strangeness-mixed matter as $K_{0}=240$ MeV, $a_{\mathrm{sym}}=32$ MeV, and $L_{\mathrm{sym}}=60$ MeV, which are considered as our final set of the parameters. We compare the current results with the observed data for the neutron stars. The solid curve designates the result for the neutron stars. When we add the hyperons or the strange quarks to neutron matter, the EoS gets softer, as discussed previously. Consequently, the mass-radius relations become very different from those of the neutron stars. The most prominent feature is that the central density starts to increase, as the strangeness content increases. Assuming that the recently found neutron stars PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] to be strangeness-mixed stars, the corresponding central densities become larger as the strangeness content increases. It also indicates that the strangeness-mixed stars become more compact than the neutron stars, which is understandable. In particular, when the central density gets lower, the mass-radius relations of the strangeness-mixed stars exhibit the propensity to be more shifted to the left. In Table 3, varying the strangeness content from 0 to 0.25, we list the corresponding results for the central densities and radii of the strangeness-mixed stars with two different masses $M=1.4M_{\odot}$ and $M=2.0M_{\odot}$ fixed.
Table 3: The central densities and radii of the strangeness-mixed stars with two different masses, $M=1.4M_{\odot}$ and $M=2.0M_{\odot}$, fixed. $n$ denotes the central number density, $\lambda$ is the corresponding ratio $n/n_{0}$, and $R$ stands for the radius of the strangeness-mixed star. We vary the strangeness content $\delta_{s}$ from 0 to 0.25.
[…]
   $\delta_{\mathrm{s}}$ $n$ $\lambda$ $R$ $n$ $\lambda$ $R$
$[\mathrm{fm}^{-3}]$ [dimensionless] [km] $[\mathrm{fm}^{-3}]$ [dimensionless] [km]
$\delta_{\mathrm{s}}=0$ 0.417 2.609 12.426 0.541 3.384 12.272
$\delta_{\mathrm{s}}=0.05$ 0.423 2.641 12.239 0.544 3.402 12.173
$\delta_{\mathrm{s}}=0.10$ 0.428 2.672 12.064 0.548 3.424 12.075
$\delta_{\mathrm{s}}=0.15$ 0.433 2.706 11.896 0.552 3.450 11.977
$\delta_{\mathrm{s}}=0.20$ 0.439 2.743 11.736 0.557 3.480 11.879
$\delta_{\mathrm{s}}=0.25$ 0.445 2.782 11.580 0.562 3.514 11.780
V Summary and outlook
We have investigated the properties of the neutron stars and strangeness-mixed stars, using the equation of the state for both neutron matter and strangeness-mixed matter, which were derived from the medium-modified pion mean-field approach (chiral quark-soliton model) [72]. All the relevant parameters were already fixed in Ref. [72]. A great virtue of this approach is that the medium modifications of the nucleon and hyperons have been considered, in addition to describing the bulk properties of nuclear media. In addition, we have considered the leptonic effects, which mainly come from the crust of the neutrons stars. The results for the mass-radius relations were well described in the current work. The central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] were obtained to be around $3.5\rho_{0}$.
We also examined the dependence of the mass-radius relations on the incompressibility $K_{0}$. As $K_{0}$ increases, central density becomes lower and the radius increases. To investigate the strangeness-mixed stars, we first examined the equation of state for strangeness-mixed matter, as the strangeness content varied from 0 to 1, which correspond to neutron matter and pure strange matter, respectively. At $\lambda\approx 0.16$, the binding energy becomes negative and turns positive, which behaves as normal nuclear matter. As the content of the strangeness increases, the binding energy becomes further lower, though strange matter has a shallow saturation point in comparison with nuclear matter. As a result, the pressure of strange-mixed matter becomes lessened as the strangeness content increases. Note that the pressure is increased by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softened. Thus, the masses of the strangeness-mixed star become smaller at the same central density. We conclude that the strange stars are more compact than the neutron stars.
Acknowledgments
The authors are grateful to Myung-Ki Cheoun for fruitful discussion and valuable comments related to leptonic effects on neutron stars. The present work was supported by Basic Science Research Program through the National Research Foundation (NRF) of Korea funded by the Korean government (Ministry of Education, Science and Technology, MEST), Grant Numbers: 2023R1A2C1008137 (Gh.-S. Y.), RS-2025-00513982(H.-Ch.K.), and 2020R1F1A1067876 (U. Y.).